# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
When $Y$ is smooth and the central point $p\in\overline{P}$ is in the interior of $\overline{P}$, then we have $\overline{X}^{v_{x}}_{\mathscr{P}}\simeq\mathbb{A}^{2}$ for every 00-dimensional stratum $x$ of $D$, and so $\overline{\mathcal{X}}_{\mathscr{P},0}$ is isomorphic to the $n$-vertex as in [41, §1.3], where $n$ is the number of irreducible components of $D$. In general, by Lemma 4.3, the singularities of the irreducible components $\overline{X}^{v_{x}}_{\mathscr{P}}$ at the central point $x_{0}$ are flat deformations of cyclic quotient singularities, and so are also cyclic quotient singularities by [68, Corollary 7.15].
Lemma 4.7.
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. Then, the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ is a quasi-projective open Kulikov surface.
Proof.
It follows from the description of the irreducible components of $\mathcal{X}_{\mathscr{P},0}$ in Lemmas 4.3 and 4.5 that $\mathcal{X}_{\mathscr{P},0}$ is an open Kulikov surface. To prove quasi-projectivity, first recall that, by Definition 2.13, the good polyhedral decomposition $\overline{\mathscr{P}}$ is regular. Hence, there exists a convex piecewise-linear function $\varphi$ on $\overline{P}$ inducing $\overline{\mathscr{P}}$. We denote by $L_{\varphi}$ the corresponding ample line bundle on $\mathcal{X}_{\overline{\mathscr{P}}}$. It follows from the toric description of the deformation in Step I of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2 that the restrictions of $L_{\varphi}$ to $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v}_{\overline{\mathscr{P}}}$ deform into ample line bundles $\widetilde{L}_{\varphi}^{v}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$. As a toric surface $X$ satisfies $H^{2}(X,\mathcal{O}_{X})=0$, it follows from deformation theory for line bundles (see for example the proof of [85, Theorem 2.5.23 (ii)]) that, in Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, the ample line bundles $L^{v}_{\varphi}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$ deform into ample line bundles $L^{v}_{\varphi}$ on the log Calabi–Yau surfaces $X^{v}_{\mathscr{P}}$. Moreover, the line bundles $L^{v}_{\varphi}$ and $L^{v^{\prime}}_{\varphi}$ are isomorphic in restriction to $X^{e}_{\mathscr{P}}$, for every edge $e\in\mathscr{P}^{[1]}$ adjacent to two vertices $v,v^{\prime}\in\mathscr{P}^{[1]}$. Hence, it follows from the description of $\mathrm{Pic}(\mathcal{X}_{\mathscr{P},0})$ in Lemma 3.10 that the ample line bundles $L^{v}_{\varphi}$ glue into a global ample line bundle on $\mathcal{X}_{\mathscr{P},0}$. ∎
4.3. Semistable mirrors of polarized log Calabi–Yau surfaces
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface with a maximal degeneration defined by $(P,\mathscr{P})$. By Theorem 3.21, up to a locally trivial deformation, we can assume that the open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$ is $d$-semistable and generic. Therefore, by Theorem 3.26, there exists a smoothing $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\to\Delta$ of $\mathcal{X}_{\mathscr{P},0}$, which in addition is an open Kulikov degeneration. From now on we refer to this smoothing as the semistable mirror to $(Y,D,L)$.
Remark 4.8.
If $(Y,D)$ is a toric pair, given by a toric surface $Y$ with its anti-canonical boundary divisor $D$, then the affinization $\overline{\mathcal{X}}_{\mathscr{P}}\rightarrow\Delta$ of the semistable mirror is the completion along $0\in\mathbb{A}^{1}$ of the three dimensional Gorenstein affine toric variety, whose fan is the cone over the momentum polytope $P$ of $(Y,D,L)$ and so which admits a natural toric morphism to $\mathbb{A}^{1}$. The semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ is the completion along 00 of the toric crepant resolution of this affine toric singularity defined by the regular triangulation $\mathscr{P}$ of $P$ into triangles of size one.
Finally, note that while a smoothing of $\mathcal{X}_{\mathscr{P},0}$ is not unique, it is unique up to deformation. In the remaining part of this paper, we will always consider smoothings of $\mathcal{X}_{\mathscr{P},0}$ up to deformation. Therefore by abuse of notation we will call $\mathcal{X}_{\mathscr{P}}$ “the” smoothing of $\mathcal{X}_{\mathscr{P},0}$, and refer to it as the semistable mirror to the polarized log Calabi–Yau surface $(Y,D,L)$, with given Symington polytope $P$ endowed with a good polyhedral decomposition $\mathscr{P}$. In the following section §5, we show how to generalize the intrinsic mirror symmetry construction of Gross–Siebert for applying it to quasi-projective open Kulikov degenerations such as $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. Eventually, we show that starting with a polarized log Calabi–Yau $(Y,D,L)$, constructing its semistable mirror, and then taking the intrinsic mirror to the semistable mirror indeed leads to a family of polarized log Calabi–Yau’s deformation equivalent to the initial $(Y,D,L)$, as shown in §8. Therefore, the semistable mirror construction can be thought of as the reverse of a generalized intrinsic mirror construction of Gross–Siebert for families of log Calabi–Yau’s. Hence, the construction of the semistable mirror provides a way to talk of the “mirror of the mirror” of a polarized log Calabi–Yau $(Y,D,L)$ with the choice of a Symington polytope $P$ and a good polyhedral decomposition $\mathscr{P}$.
Remark 4.9.
@@ 4.3. Semistable mirrors of polarized log Calabi–Yau surfaces
Let $(Y,D,L)$ be a log Calabi–Yau surface with an ample line bundle $L$. Since $L$ is ample, we have $L\cdot\beta\in\mathbb{N}$ for every effective curve class $\beta\in NE(Y)$. Hence, we have a map of monoids $NE[Y]\to\mathbb{N}$ defined by $\beta\mapsto L\cdot\beta$, inducing a formal curve $\mathrm{Spf}\,\mathbb{C}[\mathbb{N}]\to\mathrm{Spf}\,\mathbb{C}[NE(Y)]$. If $Y$ smooth, the central fiber $\overline{\mathcal{X}}_{\mathscr{P},0}$ is the $n$-vertex, obtained by gluing $\mathbb{A}^{2}$’s pairwise along $\mathbb{A}^{1}$’s as in [41]. In this case, $\overline{\mathcal{X}}_{\mathscr{P}}$ is expected to be deformation equivalent to the restriction to $\mathrm{Spf}\,\mathbb{C}[\mathbb{N}]\to\mathrm{Spf}\,\mathbb{C}[NE(Y)]$ of the mirror family $\overline{\mathcal{X}}_{GHK}\to\mathrm{Spf}\,\mathbb{C}[NE(Y)]$ to $(Y,D)$ obtained by Gross–Hacking–Keel in [41] by smoothing the $n$-vertex.
@@ 5.1. Intrinsic mirror symmetry for projective degenerations
The intrinsic mirror symmetry construction particularly applies to a projective morphism $g\colon\mathcal{X}\to S$ from a smooth variety $\mathcal{X}$ with $K_{\mathcal{X}}=0$ to a non-singular, non-complete curve $S$ with single closed point $0\in S$, such that the central fiber $\mathcal{X}_{0}=g^{-1}(0)$ is a reduced simple normal crossing divisor on $\mathcal{X}$ – this is a special situation of a so called “log smooth” morphism as in [48, §1.1, pg 7]. In particular, the schemes $\mathcal{X}$ and $S$ are viewed as “log schemes”, with a log structure defined by the data of the divisors $\mathcal{X}_{0}\subset\mathcal{X}$ and $0\in S$ respectively.
@@
as explained in [48], where $N_{pqr}^{\beta}(\mathcal{X})$ is a punctured Gromov–Witten invariant of $\mathcal{X}$, enumerating punctured log maps, that are given by rational curves of class $\beta$ with three marked points $x_{p},x_{q},x_{r}$ with contact orders $p,q,-r$ with $\mathcal{X}_{0}$. In addition, the image of $x_{r}$ is required to be a given point in the interior of the stratum $Z_{r}$ of $\mathcal{X}$ corresponding to the smallest cone of $\Sigma(\mathcal{X})$ containing $r$. Note that such curves, with negative order of tangency $-r\neq 0$ live in the log scheme $\mathcal{X}_{0}$, where the log structure is obtained by restricting the log structure on $\mathcal{X}$.
@@ 5.2. Mirror algebras
We cannot directly apply the intrinsic mirror symmetry construction of [48] to an open Kulikov degeneration $\pi\colon\mathcal{X}\to\Delta$ for three reasons. First, $\mathcal{X}$ is not an algebraic variety over an algebraic curve, but only a formal scheme over the formal disk $\Delta$. Second, the map $\pi:\mathcal{X}\rightarrow\Delta$ is not projective. Third, the central fiber $\mathcal{X}_{0}$ is a normal crossing but possibly not simple normal crossing. We will address these issues to define punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$ for $\mathcal{X}$. The first issue will be addressed by reformulating the constructions of [48] using only the special fiber $\mathcal{X}_{0}$, which is a quasi-projective variety. To resolve the second issue, we will use instead that the contraction $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is projective. Finally, the third issue is addressed in Remark 5.1.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. As $\pi$ is a smoothing of $\mathcal{X}_{0}$, there is a natural log structure on $\mathcal{X}_{0}$, log smooth over the standard log point – see e.g. [93, §3.2] for foundational aspects of log structures on formal schemes. We denote by $\mathcal{X}_{0}^{\dagger}$ the corresponding log scheme. Its tropicalization $\Sigma(\mathcal{X}_{0}^{\dagger})$ is the cone $(C(P),C(\mathscr{P}))$ over the dual intersection complex $(P,\mathscr{P})$.
We will define the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ of the mirror algebra to $\pi:\mathcal{X}\rightarrow\Delta$ using moduli spaces of stable punctured maps to $\mathcal{X}_{0}^{\dagger}$. A punctured map to $\mathcal{X}_{0}^{\dagger}$ is a log morphism $\nu^{\dagger}:C^{\dagger}\rightarrow\mathcal{X}_{0}^{\dagger}$, where $C$ is a log curve satisfying the conditions in [2, Definition 2.13]. In particular, the scheme $C$ underlying $C^{\dagger}$ is a projective curve with at worst nodal singularities, and there are distinguished points $x_{i}$ on $C$ called punctured points. Moreover, each punctured point $x_{i}$ has a contact order $p_{i}$, which is either an element of $C(P)_{\mathbb{Z}}$ or of $-C(P)_{\mathbb{Z}}$. A punctured point $x_{i}$ is called a marked point if $p_{i}\in C(P)_{\mathbb{Z}}$. Finally, a punctured map is called stable if the underlying map of schemes $\nu:C\rightarrow\mathcal{X}_{0}$ is a stable map. We refer to [2] for the details of the theory of stable punctured maps to a log scheme.
@@
The theory of punctured maps is developed in [2] under the assumption that the target has a log structure defined with respect to the Zariski topology (and not with respect to the étale topology). Correspondingly, the mirror construction of Gross–Siebert [48, 49] is formulated under the assumption that the central fiber is simple normal crossing instead of just normal crossing. If the open Kulikov surface $\mathcal{X}_{0}$ is not simple normal crossing, there is an edge $e$ in $\mathscr{P}$ connecting a vertex $v$ to itself. Taking the cone $C(\widetilde{\mathscr{P}})$ over a rational triangulation $\widetilde{\mathscr{P}}$ refining $\mathscr{P}$ and containing as vertices some rational points in the interior of every such edges $e$ defines a log blow-up $\widetilde{\mathcal{X}}$ of $\mathcal{X}$, so that the central fiber $\widetilde{\mathcal{X}}_{0}$ is now simple normal crossing, and the Gross–Siebert construction (with the modifications described below to deal with the non-compactness of $\mathcal{X}_{0}$) applies to such a degeneration $\widetilde{\mathcal{X}}\rightarrow\Delta$. Since this log blow-up is harmless for all our arguments, we will no longer mention it explicitly for the remaining of this paper.
@@
Let $r\in C(P)_{\mathbb{Z}}\setminus\{0\}$, and denote by $Z_{r}$ the stratum of $\mathcal{X}_{0}^{\dagger}$ corresponding to the smallest cone of $C(P)$ containing $r$. As in [48, §3.1], one can define a moduli space $\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)$ of punctures in $\mathcal{X}_{0}$ with contact order $-r$, which is a log stack with underlying stack $Z_{r}\times B\mathbb{G}_{m}$. As in Definition [48, 3.5], for every type $\tau$ of genus zero containing a punctured point $x_{\mathrm{out}}$ with contact order $-r$. there is a log evaluation morphism at $x_{\mathrm{out}}$
@@
By [48, Proposition 3.8], a choice of a point $z$ in the interior of $Z_{r}$ induces a log morphism
$B\mathbb{G}_{m}^{\dagger}\longrightarrow\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)\,,$
with underlying morphism the inclusion $z\times B\mathbb{G}_{m}\subset Z_{r}\times B\mathbb{G}_{m}$, where $B\mathbb{G}_{m}^{\dagger}$ is the stack $B\mathbb{G}_{m}$ endowed with the restriction of the divisorial log structure on $[\mathbb{A}^{1}/\mathbb{G}_{m}]$ defined by $[0/B\mathbb{G}_{m}]\subset[\mathbb{A}^{1}/\mathbb{G}_{m}]$. Finally, one defines
$\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z):=\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)\times_{\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)}B\mathbb{G}_{m}^{\dagger}\,,$ (5.2)
where the fiber product is taken in the category of log stacks.
Every log scheme $X$ has an associated Artin fan $\mathcal{A}_{X}$ [3, Proposition 3.1.1]. The Artin fan is a 0-dimensional log stack, which captures the combinatorial aspects of the log structure on $X$. One can define as in (5.2) a moduli space $\mathfrak{M}(\mathcal{A}_{\mathcal{X}_{0}^{\dagger}},\tau,z)$ of prestable punctured maps to the Artin fan $\mathcal{A}_{\mathcal{X}_{0}^{\dagger}}$, and there is a natural forgetful morphism
$\phi:\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)\longrightarrow\mathfrak{M}(\mathcal{A}_{\mathcal{X}_{0}^{\dagger}},\tau,z)\,.$
Denoting by $p:C\rightarrow\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ the universal domain curve, and by $\nu:C\rightarrow\mathcal{X}_{0}^{\dagger}$, then, as in [2, Proposition 3.12], $Rp_{\star}(\nu^{\star}(T_{\widetilde{X}^{\dagger}}|_{\mathcal{X}_{0}})(-x_{\mathrm{out}}))$ is a 2-term perfect obstruction theory relative to $\phi$ of relative virtual dimension zero, where $T_{\widetilde{X}^{\dagger}}|_{\mathcal{X}_{0}^{\dagger}}$ is the restriction to $\mathcal{X}_{0}$ of the log tangent bundle of $\widetilde{\mathcal{X}}^{\dagger}$.
If all the punctured points of the type $\tau$ apart from $x_{\mathrm{out}}$ are marked points, that is, with contact orders in $C(P)$, then, as in [2, Proposition 3.19], $\mathfrak{M}(\mathcal{A}_{\mathcal{X}_{0}^{\dagger}},\tau,z)$ is pure-dimensional of dimension $n-2$, and so $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ carries a virtual fundamental class
$[\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)]^{\mathrm{virt}}$ (5.3)
of dimension $n-2$, where $n$ is the number of marked points in $\tau$. In order to define invariants from these virtual fundamental classes, one needs to know that the moduli spaces $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ are proper. This properness is not immediate as $\mathcal{X}_{0}$ is not projective, but nevertheless holds by the following result. The proof of properness uses in an essential way the existence of a contraction $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ for an open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$.
@@
Let $r\in C(P)_{\mathbb{Z}}\setminus\{0\}$ and $\tau$ be a genus zero type of punctured maps to $\mathcal{X}_{0}^{\dagger}$, of class $\beta\in NE(\mathcal{X}_{0}/\overline{\mathcal{X}}_{0})\cong NE(\mathcal{X}/\overline{\mathcal{X}})$, with a punctured point $x_{\mathrm{out}}$ of contact order $-r$, and whose all other punctured points are marked points. Then, for every point $z$ in the interior of $Z_{r}$, the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ is proper.
@@
Let $\nu\colon C\to\mathcal{X}_{0}^{\dagger}$ be a punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$. As the class $\beta\in NE(\mathcal{X}_{0})$ of the image $\nu(C)$ is contained in $NE(\mathcal{X}_{0}/\overline{\mathcal{X}}_{0})$, it follows that $\nu(C)$ is contracted to a point under the map $f\colon\mathcal{X}\to\overline{\mathcal{X}}$. By definition of the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$, the image $\nu(C)$ contains the point $z$. Thus, we have $f(\nu(C))=\{f(z)\}$, and so $\nu(C)\subset f^{-1}(\{f(z)\})$. Note that for $r\neq 0$ we have $Z_{r}\subset\mathcal{X}_{0}$ and since $z\in Z_{r}$ we have $z\in\mathcal{X}_{0}$. Therefore, we have $f(z)\in\overline{\mathcal{X}}_{0}$, and so $f^{-1}(\{f(z)\})$ is actually a subvariety of $\mathcal{X}_{0}$. Since every punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ factors through the subvariety $V:=f^{-1}(\{f(z)\})$ of $\mathcal{X}_{0}$, it follows that $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)\simeq\mathcal{M}(V^{\dagger},\tau,z)$, where $V^{\dagger}$ is the log scheme obtained by restricting the log structure of $\mathcal{X}_{0}^{\dagger}$ to $V=f^{-1}(\{f(z)\})$. Since $f\colon\mathcal{X}\to\overline{\mathcal{X}}$ is projective, the fiber $V=f^{-1}(\{f(z)\})$ of $f$ is a projective subvariety of $\mathcal{X}_{0}$, and so the properness of $\mathcal{M}(V^{\dagger},\tau,z)$ follows by [2]. ∎
We can now define the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ in our setup. For every $p,q\in C(P)_{\mathbb{Z}}$, $r\in C(P)_{\mathbb{Z}}\setminus\{0\}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, we denote by $\tau_{pqr}^{\beta}$ the type of punctured maps to $\overline{\mathcal{X}}_{0}^{\dagger}$ of genus zero, class $\beta$, with two marked points with contact order $p$ and $q$ and one punctured point with contact order $-r$. As the type $\tau_{pqr}^{\beta}$ has $n=2$ marked points, the virtual fundamental class $[\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)]^{\mathrm{virt}}$ in (5.3) has dimension $n-2=0$. Moreover, by Theorem 5.2, the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ is proper and so the degree of 00-cycles on $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ is well-defined.

### v3

@@ 4.3. Semistable mirrors of polarized log Calabi–Yau surfaces
Let $(Y,D,L)$ be a log Calabi–Yau surface with an ample line bundle $L$. Since $L$ is ample, we have $L\cdot\beta\in\mathbb{N}$ for every effective curve class $\beta\in NE(Y)$. Hence, we have a map of monoids $NE[Y]\to\mathbb{N}$ defined by $\beta\mapsto L\cdot\beta$, inducing a formal curve $\mathrm{Spf}\,\mathbb{C}[\mathbb{N}]\to\mathrm{Spf}\,\mathbb{C}[NE(Y)]$. If $Y$ is smooth, then the central fiber $\mathcal{X}^{\mathrm{can}}_{\mathscr{P},0}$ is the $n$-vertex, obtained by gluing $\mathbb{A}^{2}$’s pairwise along $\mathbb{A}^{1}$’s as in [44]. In this case, $\mathcal{X}^{\mathrm{can}}_{\mathscr{P}}$ is expected to be deformation equivalent to the restriction to $\mathrm{Spf}\,\mathbb{C}[\mathbb{N}]\to\mathrm{Spf}\,\mathbb{C}[NE(Y)]$ of the mirror family $\mathcal{X}^{\mathrm{can}}_{GHK}\to\mathrm{Spf}\,\mathbb{C}[NE(Y)]$ to $(Y,D)$ obtained by Gross–Hacking–Keel in [44] by smoothing the $n$-vertex.
@@ 5.1. Intrinsic mirror symmetry for projective degenerations
The intrinsic mirror symmetry construction particularly applies to a projective morphism $g\colon\mathcal{X}\to S$ from a smooth variety $\mathcal{X}$ with $K_{\mathcal{X}}=0$ to a non-singular, non-complete curve $S$ with single closed point $0\in S$, such that the central fiber $\mathcal{X}_{0}=g^{-1}(0)$ is a reduced simple normal crossing divisor on $\mathcal{X}$ – this is a special situation of a so called “log smooth” morphism as in [51, §1.1, pg 7]. In particular, the schemes $\mathcal{X}$ and $S$ are viewed as “log schemes”, with a log structure defined by the data of the divisors $\mathcal{X}_{0}\subset\mathcal{X}$ and $0\in S$ respectively.
@@
as explained in [51], where $N_{pqr}^{\beta}(\mathcal{X})$ is a punctured Gromov–Witten invariant of $\mathcal{X}$, enumerating punctured log maps, that are given by rational curves of class $\beta$ with three marked points $x_{p},x_{q},x_{r}$ with contact orders p,q,rp,q,-r with $\mathcal{X}_{0}$. In addition, the image of $x_{r}$ is required to be a given point in the stratum $Z_{r}$ of $\mathcal{X}$ corresponding to the smallest cone of $\Sigma(\mathcal{X})$ containing $r$. Note that such curves, with negative order of tangency $r\neq 0$ live in the log scheme $\mathcal{X}_{0}$, where the log structure is obtained by restricting the log structure on $\mathcal{X}$.
@@ 5.2. Mirror algebras
We cannot directly apply the intrinsic mirror symmetry construction of [51] to an open Kulikov degeneration $\pi\colon\mathcal{X}\to\Delta$ for three reasons. First, $\mathcal{X}$ is not an algebraic variety over an algebraic curve, but only a formal scheme over the formal disk $\Delta$. Second, the map $\pi:\mathcal{X}\rightarrow\Delta$ is not projective. Third, the central fiber $\mathcal{X}_{0}$ is a normal crossing but possibly not simple normal crossing. We will address these issues to define punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$ for $\mathcal{X}$. The first issue will be addressed by reformulating the constructions of [51] using only the special fiber $\mathcal{X}_{0}$, which is a quasi-projective variety. To resolve the second issue, we will use instead that the contraction $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is projective. Finally, the third issue is addressed in Remark 5.1.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. As $\pi$ is a smoothing of $\mathcal{X}_{0}$, there is a natural log structure on $\mathcal{X}_{0}$, log smooth over the standard log point $\mathrm{pt}_{\mathbb{N}}$ – see e.g. [97, §3.2] for foundational aspects of log structures on formal schemes. We denote by $\mathcal{X}_{0}^{\dagger}$ the corresponding log scheme. Its tropicalization $\Sigma(\mathcal{X}_{0}^{\dagger})$ is the cone $(C(P),C(\mathscr{P}))$ over the dual intersection complex $(P,\mathscr{P})$.
We will define the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ of the mirror algebra to $\pi:\mathcal{X}\rightarrow\Delta$ using moduli spaces of stable punctured maps to $\mathcal{X}_{0}^{\dagger}$. A punctured map to $\mathcal{X}_{0}^{\dagger}$ is a log morphism $\nu^{\dagger}:C^{\dagger}\rightarrow\mathcal{X}_{0}^{\dagger}$, where $C^{\dagger}$ is a log curve satisfying the conditions in [2, Definition 2.13]. In particular, the scheme $C$ underlying $C^{\dagger}$ is a projective curve with at worst nodal singularities, and there are distinguished points $x_{i}$ on $C$ called punctured points. Moreover, each punctured point $x_{i}$ has a contact order $p_{i}$, which is either an element of $C(P)_{\mathbb{Z}}$ or of $C(P)_{\mathbb{Z}}$. A punctured point $x_{i}$ is called a marked point if $p_{i}\in C(P)_{\mathbb{Z}}$. Finally, a punctured map is called stable if the underlying map of schemes $\nu:C\rightarrow\mathcal{X}_{0}$ is a stable map. We refer to [2] for the details of the theory of stable punctured maps to a log scheme.
@@
[…]
@@
Let $r\in C(P)_{\mathbb{Z}}$. For $r\neq 0$, denote by $Z_{r}$ the stratum of $\mathcal{X}_{0}^{\dagger}$ corresponding to the smallest cone of $C(P)$ containing $r$. For $r=0$, set $Z_{0}:=\mathcal{X}_{0}^{\dagger}$. As in [51, §3.1], one can define a moduli space $\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)$ of punctures in $\mathcal{X}_{0}$ with contact order r-r, which is a log stack with underlying stack $Z_{r}\times B\mathbb{G}_{m}$. Following [51, Definition 3.5], for every type $\tau$ of genus zero containing a punctured point $x_{\mathrm{out}}$ with contact order r-r, there is a log evaluation morphism at $x_{\mathrm{out}}$
@@
Fix a morphism of log stacks
$W=\mathrm{pt}_{Q}/\mathbb{G}_{m}\longrightarrow\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)$ (5.2)
as in [51, §7.2] (see also [52, Step I of the proof of Theorem 6.4]), where $\mathrm{pt}_{Q}$ is a log point as in (3.2). Forgetting the log and stacky data, this is just the data of a point $z_{W}\in Z_{r}$. Then, we define
$\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}:=\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)\times_{\mathscr{P}(\mathcal{X}_{0}^{\dagger},r)}W\,$,
where the fiber product is taken in the category of fine and saturated log stacks. If all the punctured points of the type $\tau$ apart from $x_{\mathrm{out}}$ are marked points, that is, with contact orders in $C(P)$, then, as in [2, Definition 7.9], the smoothness of $\mathcal{X}$ over $\mathbb{C}$ implies that $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ carries a virtual fundamental class
$\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}]^{\mathrm{virt}}$ (5.3)
of dimension n2n-2, where $n$ is the number of marked points in $\tau$. To define invariants from these virtual fundamental classes, it is essential to establish that the moduli spaces $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ are proper. This properness is not immediate since $\mathcal{X}_{0}$ is not projective, but nevertheless holds by the following result. The proof relies on the existence of a projective contraction $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ for an open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Let $r\in C(P)_{\mathbb{Z}}$ and $\tau$ be a genus zero type of punctured maps to $\mathcal{X}_{0}^{\dagger}$, of class $\beta\in NE(\mathcal{X}_{0}/\mathcal{X}^{\mathrm{can}}_{0})\cong NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, with a punctured point $x_{\mathrm{out}}$ of contact order r-r, and whose all other punctured points are marked points. Then, for every $W$ as in (5.2), the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ is proper.
@@
Let $\nu\colon C\to\mathcal{X}_{0}^{\dagger}$ be a punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$. By definition of $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$, the image curve $\nu(C)$ contains the point $z_{W}\in Z_{r}\subset\mathcal{X}_{0}$. Thus, denoting by $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ the affinization morphism, we have $f(\nu(C))=\{f(z_{W})\}$, and so $\nu(C)\subset f^{-1}(\{f(z_{W})\})$. Since every punctured map defining a point in the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ factors through the subvariety $V:=f^{-1}(\{f(z_{W})\})$ of $\mathcal{X}_{0}$, it follows that $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}\simeq\mathcal{M}(V^{\dagger},\tau)_{W}$, where $V^{\dagger}$ is the log scheme obtained by restricting the log structure of $\mathcal{X}_{0}^{\dagger}$ to $V$. Since $f\colon\mathcal{X}\to\mathcal{X}^{\mathrm{can}}$ is projective, the fiber $V=f^{-1}(\{f(z_{W})\})$ of $f$ is a projective subvariety of $\mathcal{X}_{0}$, and so the properness of $\mathcal{M}(V^{\dagger},\tau)_{W}$ follows by [2, Corollary 3.19]. ∎
We can now define the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ in our setup, following [51, Definition 7.9]. For every $p,q,r\in C(P)_{\mathbb{Z}}$, and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, we denote by $\tau_{pqr}^{\beta}$ the type of punctured maps to $(\mathcal{X}^{\mathrm{can}}_{0})^{\dagger}$ of genus zero, class $\beta$, with two marked points with contact order $p$ and $q$ and one punctured point with contact order r-r. As the type $\tau_{pqr}^{\beta}$ has $n=2$ marked points, the virtual fundamental class $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}]^{\mathrm{virt}}$ in (5.3) has dimension n2=0n-2=0. Moreover, by Theorem 5.2, the moduli space $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ is proper and so the degree of 0-cycles on $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ is well-defined.